the two unit vectors parallel to the line 𝑦=−(3𝑥+5)



1
Expert's answer
2022-02-01T15:10:48-0500

We have for coordinates of the vector "\vec u:" "y=-3x." Then

"|\vec u|=\sqrt{x^2+y^2}=\sqrt{x^2+(-3x)^2}=|x|\sqrt{10}=1"

"|x|=1/\sqrt{10}"

The two unit vectors parallel to the line "\ud835\udc66=\u2212(3\ud835\udc65+5)" are


"\vec u_1=\lang -1/\sqrt{10}, 3/\sqrt{10}\rangle, \vec u_2=\lang 1/\sqrt{10}, -3/\sqrt{10}\rangle"